# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v13

@@ Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, which can be used for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experimental results demonstrate that the enhanced algorithms achieve significantly improved convergence speed compared to traditional first-order gradient methods. Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations. There is a great chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that quadratic gradient could be applied to a wide range of numerical optimization problems.
Index Terms:
Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
@@ I Introduction
    We propose a systematic framework to determine the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], and to assess the feasibility of employing fixed Hessian matrices for efficient computation.
@@ II Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
@@ II-A Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [11]. FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext [12] also contributes to the slowdown. Cheon et al. [13] proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in [3, 1, 14].
@@ II-C Logistic Regression Model
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
@@ III-A Simplified Fixed Hessian
Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $-\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
@@
which is the same as the formula of the naive gradient $ascent$ method. Such a coincidence not only helps generate the idea behind this work but also leads us to believe that there is a connection between the Hessian matrix and the learning rate of the gradient (descent) method.
We introduce $B^{-1}\cdot\nabla_{\boldsymbol{\beta}}\ell(\boldsymbol{\beta})$ as a novel enhanced gradient variant, and allocate a dedicated learning rate to better capture its dynamics. As long as we ensure that this new learning rate decreases from a positive floating-point number greater than 1 (such as 2) to 1 in a bounded number of iteration steps, the fixed Hessian Newton’s method guarantees the algorithm will converge eventually.
@@ III-B Quadratic Gradient: Definition
Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has gradient $\boldsymbol{g}$ and Hessian matrix $H$. For the $maximization$ problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\preceq H$ in the Loewner partial ordering. We represent these quantities as:
@@
where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$. Formally, $\tilde{B}$ is given by:
@@
By definition of the Loewner ordering, we have to prove the difference matrix $C=A-B$ is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative. By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$. By means of Gerschgorin’s circle theorem, we can bound every eigenvalue $\lambda$ of $C$ in the sense that $|\lambda-C_{ii}|\leq\sum_{i\neq j}|C_{ij}|$ for some index $i\in\{1,2,\ldots,n\}$. We conclude that $\lambda\geq A_{ii}+\epsilon+|A_{ii}|\geq\epsilon>0$ for all eigenvalues $\lambda$ and thus that $B\leq A$. ∎
@@
Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$. In this context, we need to construct $\tilde{B}$ using either a good $lower$ bound $\bar{H}$ for the Hessian $-H$ of $-F(x)$ or a good $upper$ bound $\bar{H}$ for the Hessian $H$ of $F(x)$. It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
Böhning and Lindsay [10] did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. In this work, we present a systematic approach for identifying if there were fixed Hessian matrices existing. Our proposed method serves as one of many possible approaches to detect the existence of a fixed Hessian matrix. Specifically, if our approach identifies a fixed Hessian, this provides sufficient evidence for the existence of such a matrix. However, failure to detect a fixed Hessian via our solution does not imply its non-existence. In other words, detection by our technique constitutes a sufficient but not necessary condition for the existence of a fixed Hessian matrix.
@@
Specifically, our method proceeds in the following steps:
  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directly from the Hessian of the given objective function.
@@
  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation. While this replacement satisfies theoretical convergence conditions, it may not always yield a tight bound to the true Hessian, potentially degrading the convergence rate in practice.
This search method can be applied to obtain the same fixed Hessian matrix $-\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
@@
3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
@@
6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
@@ III-C Quadratic Gradient: Algorithms
    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient $ascent$ method for NAG are as follows:
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We conduct a comprehensive evaluation of six optimization algorithms—namely, NAG, Adagrad, Adam, and their respective quadratic-gradient enhanced variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam)—on the benchmark datasets utilized by [3], including the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), NHANES III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from [14]: a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by [3], the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The Enhanced Adagrad algorithm employs a modified learning rate of $N_{t}=0.06+0.01$, in contrast to the default setting $N_{t}=0.01$ used by the original Adagrad algorithm. Similarly, the Enhanced Adam algorithm adopts an adjusted learning rate configuration with $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$, whereas the standard Adam algorithm follows its default settings of $\alpha=0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. For further details regarding these parameters, refer to [17]. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
@@
An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values. When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge. For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate, such as one of the form $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration index, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive scalar less than $1$ that governs the rate of decay.
@@ IV Secure Training
Both Adagrad and Adam algorithms are not practical solutions for homomorphic logistic regression due to their frequent reliance on inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Since data owner has to prepare the dataset and normalize it, it would also be practicable for the data owner to calculate the $\bar{B}$ owing to no leaking of sensitive data information.
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [3], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
@@
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
@@ V Experiments
The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32{,}768$, yielding a security level of $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to [3] for details on these parameters. Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm. Nevertheless, even with just four iterations, the proposed method achieves results that remain comparable to those of the baseline. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for encryption and evaluation. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.

### v14

@@ Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce an efficient gradient variant, called quadraticquadratic gradientgradient, which can be used for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experimental results demonstrate that the enhanced algorithms achieve significantly improved convergence speed compared to traditional first-order gradient methods. Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations. The proposed quadratic gradient approach may offer a unified framework that combines the advantages of first-order gradient descent/ascent algorithms and second-order Newton–Raphson methods, potentially extending its applicability to diverse numerical optimization tasks.
@@ I Introduction
    We introduce a systematic framework that not only identifies the existence of constant Hessian approximations in fixed-Hessian methods [10, 2], but also evaluates the practicality of employing fixed Hessian matrices for computational efficiency.
@@ II Preliminaries
We use the square brackets “[][\ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
@@ II-A Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation [11]. FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext [12] also contributes to the slowdown. Cheon et al. [13] proposed a method to construct an HE scheme with a rescalingrescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in [3, 1, 14].
@@ II-C Logistic Regression Model
where $S$ is a diagonal matrix with entries $S_{[i][i]}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
@@ III-A Simplified Fixed Hessian
Consider a special situation: if all the elements $b_{00},\ldots,b_{dd}$ had the same value $\eta$ with $\eta>0$, the iterative formula of the SFH method could be given as:
@@
which is the same as the formula of the naive gradient ascentascent method. This coincidence sparked the idea of unifying fixed-Hessian approaches with gradient-based optimization, forming the foundation of this study.
We introduce $B^{-1}\cdot\nabla_{\boldsymbol{\beta}}\ell(\boldsymbol{\beta})$ as a novel enhanced gradient variant, and allocate a dedicated learning rate to better capture its dynamics. If the new learning rate decreases from a positive value greater than one (e.g., 2) to one within a finite number of iterations, the fixed-Hessian Newton’s method theoretically guarantees that our proposed approach with this new learning-rate scheme will eventually converge.
@@ III-B Quadratic Gradient: Definition
Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has gradient $\boldsymbol{g}$ and Hessian matrix $H$. For the maximizationmaximization problem, let $\bar{H}$ be any matrix satisfying $\bar{H}\leq H$ in the Loewner partial ordering. We represent these quantities as:
@@
where $\nabla_{ij}^{2}=\nabla_{ji}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a new diagonal matrix $\tilde{B}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as $\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|$. Formally, $\tilde{B}$ is given by:
@@
By definition of the Loewner ordering, we have to prove the difference matrix C=ABC=A-B is non-negative definite, which means that all the eigenvalues of $C$ need to be non-negative. By construction of $C$ we have that $C_{ij}=A_{ij}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$ and $C_{ij}=A_{ij}$ for $i\neq j$. By means of Gerschgorin’s circle theorem, we can bound every eigenvalue $\lambda$ of $C$ in the sense that $|\lambda-C_{ii}|\leq\sum_{i\neq j}|C_{ij}|$ for some index $i\in\{1,2,\ldots,n\}$. We conclude that $\lambda\geq A_{ii}+\epsilon+|A_{ii}|\geq\epsilon>0$ for all eigenvalues $\lambda$ and thus that $B\leq A$. ∎
@@
Minimizing the function $F(x)$ is equivalent to maximizing the function F(x)-F(x). In this context, we need to construct $\tilde{B}$ using either a good lowerlower bound $\bar{H}$ for the Hessian H-H of F(x)-F(x) or a good upperupper bound $\bar{H}$ for the Hessian $H$ of $F(x)$. It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
Böhning and Lindsay [10] did not provide a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. In this work, we introduce a structured procedure for identifying if there were fixed Hessian matrices existing. Our proposed method serves as one of many possible approaches to detect the existence of a fixed Hessian matrix. Specifically, if our approach identifies a fixed Hessian, this provides sufficient evidence for the existence of such a matrix. However, failure to detect a fixed Hessian via our solution does not imply its non-existence. In other words, detection by our technique constitutes a sufficient but not necessary condition for the existence of a fixed Hessian matrix.
@@
Our method proceeds in the following steps:
  • Step 1: Construct $\tilde{B}$, an approximation of the Hessian, directlydirectly from the Hessian of the given objective function.
@@
  • Step 3: Evaluate the performance of the resulting fixed Hessian approximation derived in the previous step. While this replacement satisfies theoretical convergence conditions, it may not always yield a tight bound to the true Hessian, potentially degrading the convergence rate in practice.
This search method can be applied to obtain the same fixed Hessian matrix $\frac{1}{4}X^{\top}X$ as originally proposed by the authors in [10], in the context of binary logistic regression.
@@
3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ # $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ # $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
@@
6:  # $\epsilon$ is a small positive constant such as 1e81e-8
@@ III-C Quadratic Gradient: Algorithms
    NAG is a different variant of the momentum method to give the momentum term much more prescience. The iterative formulas of the gradient ascentascent method for NAG are as follows:
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\boldsymbol{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We conduct a comprehensive evaluation of six optimization algorithms—namely, NAG, Adagrad, Adam, and their respective quadratic-gradient enhanced variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam)—on the benchmark datasets utilized by [3], including the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), NHANES III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from [14]: a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by [3], the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The Enhanced Adagrad algorithm employs a modified learning rate of $N_{t}=0.06+0.01$, in contrast to the default setting $N_{t}=0.01$ used by the original Adagrad algorithm. Similarly, the Enhanced Adam algorithm adopts an adjusted learning rate configuration with $\alpha=0.01+0.001$ (the new learning rate), $\beta_{1}=0.9$, and $\beta_{2}=0.999$, whereas the standard Adam algorithm follows its default settings of $\alpha=0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. For further details regarding these parameters, see [17]. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\boldsymbol{\beta})$.
@@
An important observation is that, under a fixed learning rate, the enhanced algorithms achieve their better performance when the learning rate is between $1$ and $2$. By contrast, when the learning rate exceeds $3$, the algorithm almost never converges. For quadratic-gradient NAG algorithms, we recommend employing an exponentially decaying learning rate.
@@ IV Secure Training
Both Adagrad and Adam algorithms are not practical solutions for homomorphic logistic regression due to their frequent reliance on inversion operations. It seems plausible that the enhanced NAG is likely the reasonable choice for homomorphic LR training. We adopt the enhanced NAG method to implement secure logistic regression training based on HE. The difficulty in applying the quadratic gradient is to invert the diagonal matrix $\tilde{B}$ in order to obtain $\bar{B}$. We leave the computation of matrix $\bar{B}$ to data owner and let the data owner upload the ciphertext encrypting the $\bar{B}$ to the cloud. Given that the data owner prepares and normalizes the dataset, computing $\bar{B}$ on their side is feasible and secure, as it does not reveal any sensitive data.
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [3], utilizing the least squares approach to approximate the sigmoid function over the interval [8,8][-8,8]: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$ .
@@
where $\bar{B}_{[i][i]}$ represents the diagonal entry of $\bar{B}$, formed from $\frac{1}{4}X^{\top}X$.
@@ V Experiments
The parameters for HEAAN are configured identically to those in [3], with $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to 32,76832{,}768, yielding a security level of $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix $\bar{B}$} for preserving the precision of $\bar{B}$}. These parameters are described in greater detail in [3]. Since our enhanced NAG method requires one additional ciphertext multiplication compared to the baseline approach—resulting in greater modulus consumption—our scheme is limited to performing only four iterations of the enhanced NAG algorithm. Nevertheless, even with just four iterations, the proposed method achieves results that remain comparable to those of the baseline. Our learning time measurement reflects the time required for 4 iterations of our enhanced algorithms while the baseline’s reported average running time includes time for encryption and evaluation. It is worth noting that we relied on their open-source code to implement our homomorphic logistic regression training. Therefore, theoretically, the running time for one iteration of our algorithm should be very similar to theirs.